A parallel film of sodium light of wavelength

5880Å is incident on a thin glass plate of μ=1.5 such that

the angle of refraction in the plate is 60o. Calculate the

smallest thickness of the plate which will make it appear

dark at reflection.




1
Expert's answer
2021-10-13T15:40:45-0400

"2d\cos\beta=\frac{\lambda}n,\implies"

"d=\frac{\lambda}{2n\cos\beta},"

"d=0.392~\mu\text{m}."